# Arrow-Debreu Meets Kyle: Price Discovery for Derivatives

### v26

@@ 5.2 The Price Impact Term
According to the expression of $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ given by Equation (17), the price impact of $v$ on security $x_{i}$ conditional on $\omega$ is
@@
This is the covariance between the payoff profile $\eta(x_{i}|\,\cdot\,)$ and the projection profile $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})$ across signals, under the market maker’s posterior conditional on $\omega$. The expected price impact of $v$ on $x_{i}$ is given by the expected value $\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\,\cdot\,]$ of (19).
If $v$ consists of one share of security $x_{j}$, then
$\Pi_{insider}(v,\,\cdot\,;\widetilde{W})=\frac{\widetilde{W}(x_{j}|\,\cdot\,)}{\sigma_{j}^{2}}.$
Substituting this expression into (17) gives the following characterization of the cross price impact of $x_{j}$ on $x_{i}$.
@@
The intuition behind the cross price impact characterization (20) is straightforward: Demand for $x_{j}$ increases the price of $x_{i}$ if the market maker believes that a high demand for $x_{j}$ implies a high payoff of $x_{i}$—in other words, if he believes there is positive correlation between the insider’s demand $\widetilde{W}(x_{j}|\,\cdot\,)$ for $x_{j}$ and the payoff $\eta(x_{i}|\,\cdot\,)$ of $x_{i}$ across signals, and vice versa. More noise trades for $x_{j}$ reduces this price impact.
@@
Suppose a security $x_{j}$ only pays off under a signal $s_{k_{j}}$ and has zero payoff under any other signal. Suppose the market maker believes reasonably that insider demand for $x_{j}$ is positive conditional on $s_{k_{j}}$ and zero otherwise. Consider the following two scenarios:
@@
    (Positive Price Impact) Suppose security $x_{i}$ also pays off only under $s_{k_{j}}$, the same signal as $x_{j}$. Then $x_{j}$ has a positive price impact on $x_{i}$: Marginal demand for $x_{j}$ supports the market maker’s belief regarding $s_{k_{j}}$, which suggests to him that $x_{i}$ is also likely to pay off, leading him to raise $x_{i}$’s price. Indeed, the payoff $\eta(x|\,\cdot\,)$ and $\widetilde{W}(x_{j}|\,\cdot\,)$ are positively correlated, being positive at $s_{k_{j}}$ and zero otherwise. Therefore, by Corollary 5.2, $\frac{\partial}{\partial W(x_{j})}P(x_{i},\omega\kern 0.44998pt;\widetilde{W})>0$.
@@
    (Negative Price Impact) Suppose security $x_{i}$ only pays off under a different signal $s_{k}\!\neq\!s_{k_{j}}$. By Corollary 5.2, $x_{j}$ has a negative price impact on $x_{i}$. The intuition is the flip side of scenario (i).
@@ 5.3 No-Arbitrage
No-arbitrage is a general necessary condition for the existence of equilibrium. We now state this condition for our setting. Here, again, the projection profile $\Pi_{insider}$ between the insider’s trades and the market maker’s belief naturally occurs. We define the zero price impact portfolios under market maker belief $\widetilde{W}$ as those in the set $\mathcal{V}_{0}(\widetilde{W})\!\equiv\!\{W\colon\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0\}$. Intuitively, a portfolio in $\mathcal{V}_{0}(\widetilde{W})$ causes zero price impact because it is orthogonal to the market maker’s belief regarding the insider’s trading strategy. Such a portfolio must yield a zero payoff for the insider to prevent arbitrage. This is Theorem 5.4.
@@
Under market maker belief $\,\widetilde{W}$, an optimal portfolio for the insider can only exist if the zero price impact portfolios in $\mathcal{V}_{0}(\widetilde{W})$ yield zero payoff. Otherwise, the insider can obtain unbounded utility.
@@ 6 Equilibrium
(w.l.o.g.)   The market maker has the uniform prior on $S$.
Assumption 1 is made without loss of generality. Any other prior $\pi_{0}$ can be reinterpreted as if it were the equivalent case of the uniform prior $\pi_{u}$ and modified security payoffs $\{\frac{\pi_{0}(s_{k})}{\pi_{u}(s_{k})}\eta(\,\cdot\,|s_{k})\}_{s_{k}\in S}$. [^15] Our proofs go through for any payoff function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the modified payoffs. Once equilibrium is established under the uniform prior, this equivalence yields the comparative statics for varying priors. For example, if two securities both have higher payoffs conditional on $s_{k}$ relative to the other signals, their cross price impact increases with respect to the prior probability of $s_{k}$. This generalizes from the simplified setting of Section 2 (see footnote 8).
@@ 6.1 Canonical Game
In equilibrium, the insider’s zero payoff portfolios must coincide with the zero price impact portfolios. Zero payoff implies zero price impact because the market maker’s equilibrium belief is correct per Definition 1. Conversely, zero price impact implies zero payoff by the no-arbitrage condition in Theorem 5.4.
The zero payoff portfolios are those orthogonal to the linear span of $\{\eta(\,\cdot\,,s_{k})\}_{k}$—any portfolio $W$ with the property that $\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})=0$ for all $s_{k}$ gives the insider zero payoff. In other words, the insider’s payoff-relevant portfolios lie within the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$. We state this result as Proposition 6.1(i). On the other hand, the market maker’s noise-adjusted equilibrium beliefs must contain this span; otherwise, zero price impact allows arbitrage by the insider. This is Proposition 6.1(ii).
@@
(ii) Conversely, each $\eta(\,\cdot\,|s_{k})$ must lie in the linear span of $\{\frac{W^{*}(\,\cdot\,|s_{l})}{\sigma_{\,\cdot\,}^{2}}\}_{l}$ to prevent arbitrage.
To make the reduction of Proposition 6.1 explicit, for a portfolio $W(\,\cdot\,)=\sum\limits_{k=1}^{K}d_{k}\eta(\,\cdot\,|s_{k})$ in the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$, form the vector $d=(d_{k})\in\mathbb{R}^{K}$ of its expansion coefficients. Similarly, for market maker belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$ in the same linear span,
@@
form the matrix $\widetilde{D}=[\tilde{d}^{(1)}\;\cdots\;\tilde{d}^{(K)}]\in\mathbb{R}^{K\times K}$.
@@
${\bf L}$ is the contingent claims generalization of the noise-adjusted variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ from the single-asset setting (6). A larger $\bf{L}$ means greater noise-adjusted (co-)variations in payoffs across securities and signals—that is, higher information intensity. [^17] Here, ${\bf L}$ is ordered based on positive-semidefiniteness. This is a partial order. In contrast to the single-asset setting, the information intensity across contingent claims and signals is not always comparable between two specifications.
By Proposition 6.1, the columns of ${\bf L}$ span the expansion coefficients of the insider’s payoff-relevant portfolios in equilibrium. Therefore, we can impose the transformation
@@
Let $\widehat{D}$ denote the market maker’s belief $\widetilde{D}$ under this transformation, i.e.,
@@
Remark 6.2.
In equilibrium, the columns of ${\bf L}$ span the expected values of the sufficient statistic $\Pi_{mm}(\omega,\,\cdot\,;\widetilde{W})$. The transformation (22) normalizes the game by the insider’s information intensity. For example, when the $\eta(\,\cdot\,|s_{k})$’s have disjoint supports on $X$, this normalization simply rescales each $\eta(\,\cdot\,|s_{k})$ by $\frac{1}{c_{k}}$ where $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$. This normalization yields an isomorphic game that is symmetric across signals and invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$.
@@
Under Assumption 1, the Bayesian trading game between the insider and market maker is isomorphic to the canonical game, which is a pseudo-trading game defined as follows:
@@
  • Conditional on signal $s_{k}$, the pseudo-security $k$ has payoff of $1$ and the others have payoffs of 00.
  • The insider observes the signal. The market maker has uniform prior on signals.
  • The insider submits orders $\hat{d}\in\mathbb{R}^{K}$ for the pseudo-securities. The market maker receives order flow $\hat{\omega}=\hat{d}+\widehat{N}$, where the noise trades $\widehat{N}_{k}\!\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\!\mathcal{N}(0,1)$, $k=1,\cdots,K$.
  • The market maker has a belief $\widehat{D}$ as in (23). After receiving $\hat{\omega}$, his posterior probability for each signal $s_{k}$ (the price of pseudo-security $k$) is
@@
  • Conditional on $s_{k}$ and given market maker belief $\widehat{D}$, the insider maximizes his expected profit
@@
    where $\hat{\pi}_{1}(\hat{d}\kern 0.44998pt;\widehat{D})$ is the expectation of the pseudo-security prices (24) over the distribution of $\hat{\omega}$.
@@
An equilibrium in the canonical game is specified by a strategy matrix $D^{*}=[\delta^{(1)}\;\cdots\;\delta^{(K)}]\in\mathbb{R}^{K\times K}$ such that the strategy $\delta^{(k)}\in\mathbb{R}^{K}$ solves the insider’s problem (25) conditional on each $s_{k}$ and given market maker belief $D^{*}$, i.e.,
@@
Under the isomorphism between the two games, an equilibrium $D^{*}$ of the canonical game maps to the equilibrium of the original trading game (see Definition 1) specified by
@@
When the $\eta(\,\cdot\,|s_{k})$’s are linearly dependent, ${\bf L}^{-1}$ denotes the pseudoinverse of ${\bf L}$ (as slight abuse of notation). Compared to the case of linearly independent $\eta(\,\cdot\,|s_{k})$’s, the only caveat is that the informed demand portfolio $W^{*}(\,\cdot\,|s_{k})$ of (27) is unique only up to its payoff but not the allocations of its constituent $\eta(\,\cdot\,|s_{l})$’s.
@@
Let $p\in[0,1]^{I}$ denote the random vector (24) of the market maker’s posterior in the canonical game (with the dependence on $\hat{\omega}$ and $\widehat{D}$ understood) and $\mbox{diag}(p)\in\mathbb{R}^{K\times K}$ denote the matrix with $p$ as its diagonal and zeros elsewhere. Then, conditional on observing $s_{k}$ and given market maker belief $\widehat{D}$, the first-order condition for the insider’s problem (25) in the canonical game is
@@
where $e_{k}$ is the $k$-th standard basis vector, and the expectation $\mathbb{E}^{(\hat{d};\widehat{D})}[\,\cdot\,]$ is taken over the distribution of $\hat{\omega}$ induced by insider choice $\hat{d}$ under market maker belief $\widehat{D}$. The terms in the first-order condition (28) are the isomorphic counterparts to those in Theorem 5.1.
@@ 6.2 Symmetric Equilibrium
As noted in Remark 6.2, the canonical game is symmetric—up to permutation on $k$, the game is identical conditional on each signal $s_{k}$. This suggests an equilibrium structure of the form
@@
where ${\bf Q}=\mathbf{I}-\frac{1}{K}\bar{e}\bar{e}^{T}\in\mathbb{R}^{K\times K}$ with $\mathbf{I}$ being the identity matrix and $\bar{e}$ is the vector of $1$’s.
In other words, conditional on $s_{k}$, the insider buys $\alpha(1-\frac{1}{K})$ shares of pseudo-security $k$ and sells the other pseudo-securities for $\frac{\alpha}{K}$ shares each. This demand corresponds to the $k$-th column of the $D^{*}$ in (29). By substituting this equilibrium postulate into the first-order condition (28) conditional on each signal $s_{k}$ and collating the resulting vector equations side-by-side, we derive a matrix equation
@@
for some $\Phi\colon[0,\infty)\rightarrow\mathbb{R}$. We prove that each first-order condition is sufficient for optimality at such a fixed point of the collated first-order conditions (30). Therefore, an equilibrium is given by a solution $\alpha^{*}>0$ to the equation $\Phi(\alpha)=0$. This is Theorem 6.6.
@@
There exists $\alpha^{*}>0$ such that $\alpha^{*}{\bf Q}$ is an equilibrium of the canonical game. Equivalently, $\alpha^{*}{\bf L}^{-1}{\bf Q}$ is an equilibrium of the original trading game specified in the canonical form of Corollary 6.4.
@@
(Theorem 6.6, $K=2$)
Consider the case where $S=\{s_{1},s_{2}\}$. Substituting the equilibrium postulate (29) $D^{*}=\frac{1}{2}\begin{bmatrix}\alpha&-\alpha\\ -\alpha&\alpha\end{bmatrix}$ into the insider’s first-order condition (28) conditional on, say, $s_{1}$ gives
@@
By (32), the quantities $\mathbb{E}[p_{1}]$ and $\mathbb{E}[p_{1}p_{2}]$ in (31) are moments of a logit-normal distribution and functions of $\alpha$. Write $\mathbb{E}[p_{1}]=\phi_{1}(\alpha)$ and $\mathbb{E}[p_{1}p_{2}]=\phi_{2}(\alpha)$. Define $\Phi(\alpha)=1-\phi_{1}(\alpha)-\alpha^{2}\phi_{2}(\alpha)$. Then the first-order conditions (31) conditional on $s_{1}$ and $s_{2}$ become, respectively,
@@
The endogenous constant $\alpha^{*}$ characterizing equilibrium is the insider’s optimal portfolio size in the canonical pseudo-trading game. The intuition for $\alpha^{*}$ extends from the simplified setting of Section 2—see again Figure 1b. This is demonstrated by Example 6.8.
@@
Let $K=2$ as in Example 6.7. Consider the insider’s utility and the market maker’s posterior (32) as $\alpha$ varies.
@@
$\alpha^{*}$ is determined by only the number of signals $K$—it is independent of $\eta(\,\cdot\,|\,\cdot\,)$ or $(\sigma_{i})$. $\alpha^{*}(K)$ is an increasing concave function of $K$, as shown in Figure 2. An increase in the number of signals gives the insider a more significant information advantage over the market maker, which leads to larger insider trades. However, larger insider trades lead to greater variation in informed demand across signals, diminishing the marginal gain in the insider’s information advantage.

### v27

@@ 5.2 The Price Impact Term
From the expression of $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ in (17), it follows that the price impact of the portfolio $v$ on security $x_{i}$, conditional on $\omega$, is
@@
This is the covariance between the payoff profile $\eta(x_{i}|\,\cdot\,)$ and the projection coefficient profile $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})$ across signals, under the market maker’s posterior. The expected price impact of $v$ on $x_{i}$ is given by the expected value of (19) under $\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\,\cdot\,]$.
When the portfolio $v$ consists of one share of security $x_{j}$, $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})$ is equal to $\frac{\widetilde{W}(x_{j}|\,\cdot\,)}{\sigma_{j}^{2}}.$ Substituting this into (19) yields the following characterization of the cross price impact of $x_{j}$ on $x_{i}$.
@@
The intuition underlying the cross price impact characterization in Equation (20) is straightforward: demand for $x_{j}$ increases the price of $x_{i}$ if the market maker believes that a high demand for $x_{j}$ correlates with a high payoff of $x_{i}$—in other words, if he believes there is positive correlation between the insider’s demand $\widetilde{W}(x_{j}|\,\cdot\,)$ for $x_{j}$ and the payoff $\eta(x_{i}|\,\cdot\,)$ of $x_{i}$ across signals, and vice versa. A greater noise trading intensity $\sigma_{j}$ for $x_{j}$ reduces this price impact.
@@
Suppose a security $x_{j}$ pays off only under signal $s_{k_{j}}$ and not any other signal. Suppose further that the market maker reasonably believes that the insider’s demand for $x_{j}$ is positive under $s_{k_{j}}$ but zero otherwise. Consider the following two scenarios for another security $x_{i}$:
@@
    (Positive Price Impact) Suppose $x_{i}$ also pays off only under $s_{k_{j}}$, the same signal that triggers payoff for $x_{j}$. Then $x_{j}$ has a positive price impact on $x_{i}$: a strong demand for $x_{j}$ suggests to the market maker that $x_{i}$ is also likely to pay off, leading him to raise the price of $x_{i}$. Indeed, here the payoff $\eta(x_{i}|\,\cdot\,)$ and $\widetilde{W}(x_{j}|\,\cdot\,)$ are positively correlated, both being positive at $s_{k_{j}}$ and zero otherwise. Therefore, by Corollary 5.2, $\frac{\partial}{\partial W(x_{j})}P(x_{i},\omega\kern 0.44998pt;\widetilde{W})>0$.
@@
    (Negative Price Impact) Suppose $x_{i}$ only pays off under a signal different from $s_{k_{j}}$. By Corollary 5.2, $x_{j}$ has a negative price impact on $x_{i}$. The intuition is the flip side of scenario (i).
@@ 5.3 No-Arbitrage
No-arbitrage is a general necessary condition for the existence of equilibrium. We now state this condition for our setting. Here, once again, the projection coefficient profile $\Pi_{insider}$ between the insider’s trades and the market maker’s belief naturally occurs. Given a market maker belief $\widetilde{W}$, we define the corresponding zero price impact portfolios as portfolios $W$ for which $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0$. Intuitively, such a portfolio causes zero price impact because it is orthogonal to the market maker’s belief about the insider’s trading strategy. Such a portfolio must yield a zero payoff for the insider; otherwise, arbitrage opportunities arise for the insider. This is Theorem 5.4.
@@
Given any market maker belief, an optimal portfolio for the insider can exist only if the zero price impact portfolios corresponding to that belief yield zero payoff for the insider. Otherwise, the insider can obtain unbounded utility.
@@ 6 Equilibrium
(w.l.o.g.)   The market maker has a uniform prior on $S$.
Assumption 1 is made without loss of generality. Any other prior $\pi_{0}$ can be reinterpreted as if it were the equivalent case of the uniform prior $\pi_{u}$ and modified security payoffs $\{\frac{\pi_{0}(s_{k})}{\pi_{u}(s_{k})}\eta(\,\cdot\,|s_{k})\}_{s_{k}\in S}$. [^14] Our proofs go through for any payoff function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the modified payoffs. Once equilibrium is established under the uniform prior, this equivalence yields the comparative statics for varying the prior. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. This generalizes from the simplified setting in Section 2 (see footnote 8).
@@ 6.1 Canonical Game
In equilibrium, the insider’s zero payoff portfolios must coincide with the zero price impact portfolios. Zero payoff implies zero price impact because the market maker’s equilibrium belief is correct per Definition 1. Conversely, by the no-arbitrage condition of Theorem 5.4, zero price impact implies zero payoff.
The zero payoff portfolios are those orthogonal to the linear span of $\{\eta(\,\cdot\,,s_{k})\}_{k}$—any portfolio $W$ satisfying $\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})=0$ for all $s_{k}$ yields zero payoff for the insider. Thus, the insider’s payoff-relevant portfolios lie within the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$. We state this result as Proposition 6.1(i). On the other hand, the market maker’s noise-adjusted equilibrium beliefs must contain this span; otherwise, zero price impact allows arbitrage by the insider. This is Proposition 6.1(ii).
@@
(ii) Conversely, each $\eta(\,\cdot\,|s_{k})$ must lie in the linear span of $\{\frac{W^{*}(\,\cdot\,|s_{l})}{\sigma_{\,\cdot\,}^{2}}\}_{l}$ to prevent arbitrage opportunities.
To make the reduction of Proposition 6.1 explicit, represent a portfolio $W(\,\cdot\,)=\sum\limits_{k=1}^{K}d_{k}\eta(\,\cdot\,|s_{k})$ in the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$ by the vector $d=(d_{k})\in\mathbb{R}^{K}$ of its expansion coefficients. Similarly, represent a market maker belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$ in the same linear span,
@@
by the matrix $\widetilde{D}=[\tilde{d}^{(1)}\;\cdots\;\tilde{d}^{(K)}]\in\mathbb{R}^{K\times K}$.
@@
${\bf L}$ generalizes the noise-adjusted variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ from the single-asset setting (6). A larger $\bf{L}$ means greater noise-adjusted (co-)variations in payoffs across securities and signals, reflecting higher information intensity. [^16] Here, ${\bf L}$ is ordered based on positive-semidefiniteness. This is a partial order. Unlike the single-asset setting, the information intensity across securities and signals is not always directly comparable between two specifications.
By Proposition 6.1, the columns of ${\bf L}$ span the expansion coefficients of the insider’s payoff-relevant portfolios in equilibrium. Therefore, we can apply the canonical transformation
@@
Under this transformation, a market maker belief $\widetilde{D}$ transforms as
@@
Canonical Transformation: Intuition
The canonical transformation (22) normalizes the trading game by the insider’s information intensity. To see the intuition, suppose the $\eta(\,\cdot\,|s_{k})$’s have non-overlapping payoffs across the state space, such as in Section 2. In this case, the information intensity matrix ${\bf L}$ is diagonal, and this transformation rescales each $\eta(\cdot|s_{k})$ by $\frac{1}{c_{k}}$, where $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$ is the total noise-adjusted payoff variation across securities conditional on $s_{k}$. In turn, this transforms ${\bf L}$ to the identity matrix—thereby normalizing the information intensity across securities and signals. [^17] Here, we note again that our proofs go through for any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$ of states and signals, including the rescaled payoffs. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
@@
Under Assumption 1, the Bayesian trading game between the insider and market maker is isomorphic to the canonical game, a pseudo-trading game defined as follows:
@@
  • Conditional on signal $s_{k}$, pseudo-security $k$ has payoff of $1$ while all others have payoffs of 00.
  • The insider observes the signal, while the market maker has a uniform prior on signals.
  • The insider submits orders $\hat{d}\in\mathbb{R}^{K}$ for the pseudo-securities, and the market maker receives order flow $\hat{\omega}=\hat{d}+\widehat{N}$, where $\widehat{N}_{k}$ are i.i.d. $\mathcal{N}(0,1)$ noise trades for $k=1,\cdots,K$.
  • The market maker has a belief $\widehat{D}$ as specified in (23). Upon receiving $\hat{\omega}$, he updates the posterior probability of each signal $s_{k}$ (which is also the pseudo-security $k$ price) to
@@
  • Conditional on $s_{k}$ and given market maker belief $\widehat{D}$, the insider maximizes expected profit:
@@
    where $\hat{\pi}_{1}(\hat{d}\kern 0.44998pt;\widehat{D})$ is the expected value of the prices (24) over the distribution of $\hat{\omega}$.
@@
In the canonical game, an equilibrium is specified by a $K\times K$ strategy matrix $D^{*}=[\delta^{(1)}\;\cdots\;\delta^{(K)}]$ such that, under the same market maker belief $D^{*}$, each strategy $\delta^{(k)}\in\mathbb{R}^{K}$ solves the insider’s problem (25) conditional on $s_{k}$, i.e.,
@@
Given the isomorphism between the two games, an equilibrium $D^{*}$ of the canonical game maps to the equilibrium of the original trading game (see Definition 1) specified by
@@
When the $\eta(\,\cdot\,|s_{k})$’s are linearly dependent, ${\bf L}^{-1}$ is interpreted as the pseudoinverse of ${\bf L}$. Compared to the case of linearly independent $\eta(\,\cdot\,|s_{k})$’s, the only caveat is that the informed demand portfolio $W^{*}(\,\cdot\,|s_{k})$ of (27) is unique only up to its payoff, but not the allocations of its constituent $\eta(\,\cdot\,|s_{l})$’s.
@@
Let $p\in[0,1]^{I}$ denote the random vector (24) of the market maker’s posterior in the canonical game (with the dependence on $\hat{\omega}$ and $\widehat{D}$ understood) and $\mbox{diag}(p)\in\mathbb{R}^{K\times K}$ denote the matrix with $p$ along the diagonal and zeros elsewhere. Then, conditional on observing $s_{k}$ and given market maker belief $\widehat{D}$, the first-order condition for the insider’s problem (25) in the canonical game is
@@
where $e_{k}$ denotes the $k$-th standard basis vector, and the expectation $\mathbb{E}^{(\hat{d};\widehat{D})}[\,\cdot\,]$ is taken over the distribution of $\hat{\omega}$ induced by the insider’s choice $\hat{d}$ when the market maker holds belief $\widehat{D}$. The terms in the first-order condition (28) are the isomorphic counterparts to those in Theorem 5.1.
@@ 6.2 Symmetric Equilibrium
The canonical game is symmetric—up to permutation on $k$, the game is identical conditional on each signal $s_{k}$. This symmetry suggests an equilibrium structure of the form
@@
where ${\bf Q}=\mathbf{I}-\frac{1}{K}\bar{e}\bar{e}^{T}\in\mathbb{R}^{K\times K}$, with $\mathbf{I}$ denoting the identity matrix and $\bar{e}$ the vector of $1$’s.
This means that, conditional on $s_{k}$, the insider buys $\alpha(1-\frac{1}{K})$ shares of pseudo-security $k$ and sells the other pseudo-securities for $\frac{\alpha}{K}$ shares each. This demand corresponds to the $k$-th column of $D^{*}$ in (29). By substituting this equilibrium postulate into the first-order condition (28) conditional on each signal $s_{k}$ and collating the resulting vector equations side-by-side, we derive a matrix equation
@@
for some scalar function $\Phi\colon[0,\infty)\rightarrow\mathbb{R}$. We prove that each first-order condition is sufficient for optimality at such a fixed point of the collated first-order conditions (30). Thus, an equilibrium corresponds to a solution $\alpha^{*}>0$ to the equation $\Phi(\alpha)=0$. This is Theorem 6.5.
@@
There exists $\alpha^{*}>0$ such that $\alpha^{*}{\bf Q}$ is an equilibrium of the canonical game. Equivalently, $\alpha^{*}{\bf L}^{-1}{\bf Q}$ specifies an equilibrium of the original trading game, in the canonical form of Corollary 6.3.
@@
(Theorem 6.5, Binary Signal Case)
Suppose the signal is binary, $S=\{s_{1},s_{2}\}$. Substituting the equilibrium postulate (29) $D^{*}=\frac{1}{2}\begin{bmatrix}\alpha&-\alpha\\ -\alpha&\alpha\end{bmatrix}$ into the insider’s first-order condition (28) conditional on, say, $s_{1}$ gives
@@
By (32), the quantities $\mathbb{E}[p_{1}]$ and $\mathbb{E}[p_{1}p_{2}]$ in (31) are moments of a logit-normal distribution and functions of $\alpha$. Write $\mathbb{E}[p_{1}]$ as $\phi_{1}(\alpha)$ and $\mathbb{E}[p_{1}p_{2}]$ as $\phi_{2}(\alpha)$. Define $\Phi(\alpha)=1-\phi_{1}(\alpha)-\alpha^{2}\phi_{2}(\alpha)$. Then the first-order conditions (31) conditional on $s_{1}$ and $s_{2}$ become, respectively,
@@
The endogenous constant $\alpha^{*}$ characterizing equilibrium is the insider’s optimal portfolio size in the canonical pseudo-trading game. The intuition behind $\alpha^{*}$ builds on the simplified setting discussed in Section 2—see again Figure 1b. This is demonstrated by Example 6.7.
@@
In the same setting as Example 6.6, consider the insider’s utility and the market maker’s posterior (32), as $\alpha$ varies.
@@
$\alpha^{*}$ is determined only by the number of signals $K$—it is independent of $\eta(\,\cdot\,|\,\cdot\,)$ or $(\sigma_{i})$. $\alpha^{*}(K)$ is an increasing concave function of $K$, as shown in Figure 2. Increasing the number of signals enhances the insider’s informational advantage over the market maker, resulting in larger insider trades. However, larger insider trades also introduce greater variation in informed demand across signals, diminishing the incremental gain in the insider’s information advantage.